OpenStudy (anonymous):

u and v are column vectors in R^2 and A = I +uv^T. Show that if u^Tv is not equal to -1 then A^-1 = I -uv^T/(1+u^Tv)

OpenStudy (amistre64):

if u and v are 2 column vectors in R^2, what is uv?

OpenStudy (amistre64):

\[\binom{x_1}{y_1}\binom{x_2}{y_2}=\text{hard to do a rc operation}\]

OpenStudy (anonymous):

uv would be a scalar, right?

OpenStudy (amistre64):

matrix multipication can only be perforned if col1 = row2 right? we have 1 column on the left, and 2 rows on the right

OpenStudy (anonymous):

Your right, sorry a typo it should be u and v are column vectors in R^2 and A = I +uv^T. Show that if u^Tv is not equal to -1 then A^-1 = I -uv^T/(1+u^Tv)

OpenStudy (anonymous):

Okay I edited the question

OpenStudy (amistre64):

u = x1 v=x2 y1 y2 uT = x1 y1 vT = x2 y2 uvT = x1x2 x1y2 y1x2 y1y2 I + uvT = 1+x1x2 x1y2 = A y1x2 1+y1y2 so thats one part written out

OpenStudy (amistre64):

having trouble recalling the uv stuff. one is a dot product and the other is matrix multiplication

OpenStudy (amistre64):

uv dot product would be a scalar ... x1x2 + y1y2

OpenStudy (amistre64):

youre pretty much trying to manipulate A into A^-1 and show there respective parts. for which i believe we need a determinant for

OpenStudy (amistre64):

if you have a picture of the problem, that might clear up some stuff about the post

OpenStudy (anonymous):

Sorry there is no picture, just a word problem.

OpenStudy (amistre64):

having trouble deciphering the format of the post as given. Usually there is some notation and such given in the word problem to help distinguish some nuances

OpenStudy (amistre64):

uTv and uvT seems to relate to matrix multiplication and result in a new matrix uv would indicate a dot product that produces a scalar

OpenStudy (amistre64):

if the phrase: uTv = -1 referes to a determinant .... than it would seem apparent that 1+uTv cannot be zero ... or something

OpenStudy (amistre64):

A = 1+x1x2 x1y2 y1x2 1+y1y2 det(A) = (1+x1x2)(1+y1y2)-(x1y2)(x2y1) A^-1 = (some moving about of the components of A)/det(A)

OpenStudy (amistre64):

does the moving about of the components of A equate to 1-uvT ? and does the det(A) = (1+det(uTv)) ??

OpenStudy (amistre64):

i think the rule of a 2x2 is what, swap a d and negate b c

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

then compare the swap to: I - uvT I-uvT = 1-x1x2 -x1y2 -y1x2 1-y1y2